Anna has a vertical that is just over .5 m

1) estimate the force she exerts on the basketball court if she pushes off the ground for .3 seconds

2) compare this with the force the surface exerts on Anna

3) repeat the problem looking at the time interval when she is landing back on the floor


1
Expert's answer
2021-06-03T09:16:57-0400

1)

"P=\frac At=\frac{Fh}{t}\implies F=\frac{Pt}{h},"

"F=\frac{P\cdot 0.3}{0.5}=0.6P~\text N,"

2)

"F_1=-F=-0.6P~\text N,"

3)

"F_2=F=0.6P~\text N."